OpenStudy (anonymous):

please help, 1 equation

OpenStudy (anonymous):

What is an estimate of the solution of the equation 6n + 3 = 2? Use a table.

OpenStudy (gorv):

6n+3=2 6n=2-3 n=-1/6

OpenStudy (anonymous):

Is this okay? 6n = -1 n = -1/6

OpenStudy (gorv):

yep

OpenStudy (anonymous):

Thanks!